# The Optimization Landscape of SGD Across the Feature Learning Strength

### v1

@@ Appendix E Additional ResNet and ViT Plots
Figure 12: Multiple step-wise descent curves can be seen even in the very realistic setting of a vision transformer on TinyImageNet.
Appendix F Effect of width and check of $\mu$-consistency
@@
F.1 $\gamma$-dependency of $\mu$-consistency
@@
Figure 15: The same as 13, but for a CNN on CIFAR-5M with MSE loss.
Figure 16: The same as 14, but for a CNN on CIFAR-5M with MSE loss.
Figure 17: The same as 14, but for a CNN on CIFAR-5M with cross-entropy loss.
Figure 18: The same as 13, but for a CNN on CIFAR-5M with cross-entropy loss.
Appendix G Effect of depth
@@ G.1 MLPs
G.2 ResNets and Depth-$\mu$P
In this section we test the effect of the parameterization on deep neural networks with residual connections. Specifically, in Figure 22 and Figure 21 we evaluate the phase plots of neural networks with increasing depths parameterized in Depth-$\mu$P, as detailed in Bordelon et al. (2023); Yang et al. (2023b), and $\mu$P.
@@ G.2 ResNets and Depth-$\mu$P
Appendix H Effect of batch size
@@
Appendix I Consistency Across $\gamma$
@@ I.1 Lazy Consistency
I.2 Early Time Rich Consistency
@@
I.3 Late Time Rich Consistency
@@
Appendix J Further Hessian Plots
@@
Appendix K Further Catapult Plots
Figure 32: The catapult effect along the $\eta\sim\gamma^{2}$ line for the MSE loss for CNNs and MLPs.
@@
Appendix L Kernel-Target Alignment
@@
Appendix M Further Silent Alignment Plots
@@
Appendix N Further Function Similarities

### v3

@@ Appendix H Additional ResNet and ViT Plots
Figure 16: Stepwise loss trajectories can be seen even in the very realistic setting of a vision transformer on TinyImageNet. The trajectories here exhibit two lower plateaus in addition to the long plateau from $t=0$ until the first drop.
Appendix I Effect of width and check of $\mu$-consistency
@@
I.1 $\gamma$-dependency of $\mu$-consistency
@@
Figure 19: The same as 17, but for a CNN on CIFAR-5M with MSE loss.
Figure 20: The same as 18, but for a CNN on CIFAR-5M with MSE loss.
Figure 21: The same as 18, but for a CNN on CIFAR-5M with cross-entropy loss.
Figure 22: The same as 17, but for a CNN on CIFAR-5M with cross-entropy loss.
Appendix J Effect of depth
@@
J.2 ResNets and Depth-$\mu$P
In this section we test the effect of the parameterization on deep neural networks with residual connections. Specifically, in Figure 26 and Figure 25 we evaluate the phase plots of neural networks with increasing depths parameterized in Depth-$\mu$P, as detailed in Bordelon et al. (2023); Yang & Littwin (2023), and $\mu$P.
@@
Appendix K Effect of batch size
@@
Appendix L Consistency Across $\gamma$
@@
L.2 Early Time Rich Consistency
@@
L.3 Late Time Rich Consistency
@@
Appendix M Further Hessian Plots
@@
Appendix N Further Catapult Plots
Figure 36: The catapult effect along the $\eta\sim\gamma^{2}$ line for the MSE loss for MLPs on MNIST-1M and CIFAR-5M
@@
Appendix O Alternative Scalings of $\eta$ with $\gamma$
In this section, we briefly illustrate how taking alternative scalings of $\eta$ different from $\gamma^{2/L}$ can lead to a network that either undertrains or does not converge at large $\gamma$.
Figure 38: Sketch of phase plots with alternative scalings of $\eta$ with $\gamma$ (overlaid in dashed blue lines). We expect the $\eta\sim\gamma^{3/L}$ scaling to diverge at large $\gamma$ at any $T$ while the $\eta\sim\gamma^{2/(L+1)}$ will remain convergent, but under-train compared to the optimal learning rate
Figure 39: Alternative scalings of $\eta\sim\gamma^{3/L}$ for a) an $L=3$ MLP on MNIST-1M and b) an $L=4$ CNN on CIFAR-5M. Large $\gamma$ values not shown are those that have diverged. Indeed, for this scaling, $\gamma$ values above $10^{4}$ are found to have divergent loss dynamics. Similarly, we plot $\eta\sim\gamma^{2/(L+1)}$ in the same settings in c) and d). All of these networks train, but they do so less efficiently than the optimal scaling of $\eta\sim\gamma^{2/L}$. We illustrate the optimal scaling in figures e) and f), and observe that they achieve better final loss values than c) and d) while still allowing arbitrarily large values of $\gamma$ to train.
Appendix P Kernel-Target Alignment
@@
Appendix Q Further Silent Alignment Plots
@@
Appendix R Further Function Similarities
@@
Appendix S Weight and Activation Movement across Time and Richness
In this section, we confirm that larger values of $\gamma$ correlate directly to weight movement across a variety of norms. We consistently verify that the weights deviate as $\|W^{\ell}(t)-W^{\ell}(0)\|\sim\gamma^{1/L}$ across layers, times, and $\gamma$.
Figure 45: Weight movement $\|W^{\ell}(t)-W^{\ell}(0)\|$ for a CNN trained on CIFAR-5M. We plot these across different layers $\ell$ at different steps $t$ of training as a function of $\gamma$. We see a very sharp and clear trend that larger value of $\gamma$ correlate with meaningfully larger movement across a variety of norms. Specifically, we see that all norms of matrices scale as $\gamma^{1/L}$ after sufficient training steps. The dashed black line represents the curve $\gamma^{1/L}$ for reference. a) Frobenius norm, b) Spectral/Operator nom, c) Nuclear norm.